Your company, Jim's Whole Grains, is evaluating the potential purchase a new oven to bake sandwich buns. If purchased, the new oven will replace the existing oven, which was purchased seven years ago for $1,000,000. Depreciation on the old oven has been computed on a straight-line basis over its expected 15 year life to an ending book value of $100,000, even though you expected it to be worthless at that time. The new oven will cost $2,000,000 and will fall into the MACRS 10-year class for depreciation purchases. If you purchase it, it is expected to last for eight years, at the end of which you expect to be able to sell it for $150,000. (Note that both of the ovens, old and new, therefore have an effective remaining life of eight years at the time of your analysis.) If you purchase the new oven, you estimate that you can sell the old one for $400,000 at the same time that you purchase the new oven. The new oven has two main advantages. First, the operating costs for your current baking operations will be reduced by $125,000 per year. The second advantage is that you will be able to produce new types of buns. Sales of these new buns are expected to be $150,000 per year in the first year and grow by 3% per year throughout the life of the new oven. The associated costs of the new buns will be $60,000 in the first year and will grow by 4% each year. Since the new oven will allow you to sell these new products, you anticipate that NWC will have to increase immediately by $20,000 upon purchase of the new oven to sustain the new, higher level of operations. Your company uses a 9% cost of capital for this type of project and the marginal tax rate is 30%. Using at least two decision criteria, should this project be accepted?
Answer #1
DEPRECIATION OF OLD OVEN
A Cost of old Oven $1,000,000
B Salvage Value at end of 15 years $100,000
C=(A-B)/15 Annual Depreciation $60,000
Accumulated Depreciation today=60000*7 $420,000
Y
N Year FromToday 1 2 3 4 5 6 7 8
B1 Depreciation 0f old Oven $60,000 $60,000 $60,000 $60,000 $60,000 $60,000 $60,000 $60,000
Book Value of old oven now $580,000 (1000000-420000)
Before tax salvage value $400,000
Loss on sale $180,000
Tax Saving on loss =180000*30% $54,000
After tax Cash Flow on Salvage $454,000 (400000+54000)
DEPRECIATION OF NEW OVEN
Cost of New Oven $2,000,000
N Year(from today) 1 2 3 4 5 6 7 8
A 10 year MACRS depreciation Rate 10.00% 18.00% 14.40% 11.52% 9.22% 7.37% 6.55% 6.55%
B2=A*2000000 Depreciation $200,000 $360,000 $288,000 $230,400 $184,400 $147,400 $131,000 $131,000
Accumulated Depreciation $200,000 $560,000 $848,000 $1,078,400 $1,262,800 $1,410,200 $1,541,200 $1,672,200
Book Value at end of 8years $327,800 (2000000-1672200)
Before tax salvage value $150,000
Loss on sale $177,800
Tax Saving on loss =177800*30% $53,340
Salvage Cash flow at end of year 8 $203,340 (150000+53340)
Present value of Cash Flow=(Cash Flow)/((1+i)^N)
i=discount Rate =9%=0.09
N=Year of Cash Flow
N Year From Today 1 2 3 4 5 6 7 8
C=B2-B1 Incremental Depreciation $140,000 $300,000 $228,000 $170,400 $124,400 $87,400 $71,000 $71,000
D=C*30% Depreciation tax shield $42,000 $90,000 $68,400 $51,120 $37,320 $26,220 $21,300 $21,300
E Before tax saving in operating cost $125,000 $125,000 $125,000 $125,000 $125,000 $125,000 $125,000 $125,000
F=E*(1-0.3) After Tax Saving in Operating cost $87,500 $87,500 $87,500 $87,500 $87,500 $87,500 $87,500 $87,500
G Additional Sales Revenue from New buns $150,000 $154,500 $159,135 $163,909 $168,826 $173,891 $179,108 $184,481
H Additional Costs from New buns ($60,000) ($62,400) ($64,896) ($67,492) ($70,192) ($72,999) ($75,919) ($78,956)
I=G+H Before tax increase in Profits $90,000 $92,100 $94,239 $96,417 $98,635 $100,892 $103,189 $105,525
J=I*(1-0.3) After tax increase in Profits $63,000 $64,470 $65,967 $67,492 $69,044 $70,624 $72,232 $73,868
Terminal Cash Flow
K Incremental Salvage Cash Flow(203340-100000) $103,340
L Release of initial Net Working Capital of new Bun $20,000
M=D+F+J+K+L Total Cash Inflow $192,500 $241,970 $221,867 $206,112 $193,864 $184,344 $181,032 $306,008 SUM
K=J/(1.09^N) Present Value of Cash Inflow $176,606 $203,661 $171,322 $146,015 $125,999 $109,919 $99,031 $153,575 $1,186,127
PV Sum of Present Value of Cash inflows $1,186,127

N C=B2-B1 DEC 30% FEE*(1-0.3) Year From Today Incremental Depreciation $140,000 Depreciation tax shield $42,000 Before tax saH56 X fac =IRR(H55:P55) N O P $103,340 $20,000 $306,008 SUM $153,575 $1,186,127 $181,032 $99,031 _ H I K L M 42 Incremental S